OpenStudy (anonymous):

Solve: (u + 3)(u - 3) = 8u

OpenStudy (anonymous):

u^2 - 9 = 8u u^2 -8u -9=0 u^2 - 9u + u -9=0

OpenStudy (anonymous):

now solve it like i did

OpenStudy (anonymous):

I got u = 0; u = -3; u = 3

OpenStudy (anonymous):

?

OpenStudy (anonymous):

youll get only 2 values

OpenStudy (anonymous):

u = -1; u = 9

OpenStudy (anonymous):

now ur right

OpenStudy (anonymous):

sorry I totally mess the first one up

OpenStudy (anonymous):

Thanks !!!

OpenStudy (anonymous):

using the fomula (a+b) (a-b) = a^2 - b^2 u get u^2 - (3)^2 = 8u => u^2 - 8u - 9 = 0 now u hv to factorise -9 into two factors A and B so that AB = -9 and A + B = -8 9 * -1 = -9 and 9 + (-1) = 8 => u^2 -9u + u - 9 =0 => u(u-9) +1(u-9) = 0 => (u+1) (u-9) = 0 for the product to b zero either of the factors has to b zero if u+1 = 0 => u = -1 if u-9 = 0 => u = 9